OpenStudy (anonymous):

Kevin made contributions to a Roth IRA over the course of 40 working years. His contributions averaged $3,000 annually. Kevin was in the 25% tax bracket during his working years. The average annual rate of return on the acount was 7%. Upon retirement, Kevin stopped working and making Roth IRA contributions. What is the effective value of Kevin's Roth IRA at retirement?

OpenStudy (anonymous):

A. $ 449,179.01 B. $ 568,905.34 C. $ 580,905.34 D. $ 586,905.34

OpenStudy (tkhunny):

Have you considered accumulating the sum? Of course, it will be a very rough approximation as it is unlikely that Beginning-of-Year or End-of-Year will be useful assumptions. And that "average contribution" thing is horrible. i = 0.07 Tax = 0.25 Accumulation Interest = i*(1-Tax) = 0.07(1-0.25) = 0.07*0.75 = 0.0525 Give it a go!

OpenStudy (anonymous):

@tkhunny accumulation interest = 0.0525?

OpenStudy (tkhunny):

Accumulate away!

OpenStudy (anonymous):

...sorry but what am I accumulating? @tkhunny

OpenStudy (tkhunny):

Please reread the problem statement.

OpenStudy (anonymous):

@tkhunny I still don't follow :(

OpenStudy (tkhunny):

Why have you been given this problem? Year 1 at the end: $3000 Year 2 at the end: $3000(1.0525) + $3000 Year 3 at the end: $3000(1.0525)^2 + $3000(1.0525) + $3000 etc. Year 40 ---

OpenStudy (anonymous):

@tkhunny Is there a formula to calculate the cost of all 40 years at once without listing each year?

OpenStudy (anonymous):

Answer was B. $ 568,905.34,

OpenStudy (tkhunny):

Yes, there is. You should have seen it long before you saw this problem.